Investigate convergence or divergence of the following sequence with justification:

(√n2 + 1)(√n).



1
Expert's answer
2021-02-24T06:08:08-0500

"\sum _{n=1}^{\infty \:}\sqrt{n^2+1}\sqrt{n}"

Applying series convergence test

"\mathrm{If\:}\lim _{n\to \infty }a_n\ne 0\mathrm{\:then\:}\sum a_n\mathrm{\:diverges}"

"\lim _{n\to \infty \:}\left(\sqrt{n^2+1}\sqrt{n}\right)"

"\lim _{n\to \infty \:}\left(\left(\sqrt{n^2+1}\right)\cdot \lim _{n\to \infty \:}\left(\sqrt{n}\right)\right)"

"\infty \cdot \infty \:"

"\infty"

"\sum _{n=1}^{\infty \:}\sqrt{n^2+1}\sqrt{n}\quad \mathrm{diverges}"